2.1 Prove that for allnen, & 2+3=v* +n+3. Prov 2k + 3 = 7 n + 3. E

Use Mathematical Induction
Answer #1

- 2x+3 = n² en 3 fou dleasily À 23 - (2x) + 3 otherwise if i 24+2 = 1 4 (243) I Kai then n=2 = 12 2.5 + 3 + 2.2 +3 = 2+ 3+ 4+that is if we proof 2 EK - nen we are done. Mone for na 4 reft hand side 2.1=2 and R.H.S 1²+ 1 = 2 so for nat this is teure.that is for not it is time we assume Fan nak it is also teme and proved that fan nakt it is also tome, then by principal of